In a R-L-C AC circuit, R= 7.75 Ohm, L= 0.078 H, C=0.00016 F. Supply voltage is v= 310 sin( 410t). Calculate the power loss in R
Answer #1

I = V/Z-

Z = sqrt (R2 + (Xl - XC)2)

Xl = w*L = 410*7.8*10-2 = 31.98 ohm

XC = 1/wC = 15.24 ohm

Z = sqrt(60.06 + 280.22) = 18.44 ohm

I = 310/18.44 = 16.81

P = I2R = 2190.3 W